# The stringy geometry of integral cohomology in mirror symmetry

### v2

@@ 4 Flat gerbes and discrete torsion
In this section we turn to the situation where $X$ is simply connected but $B(X)=\{H^{3}(X,{\mathbb{Z}})\}_{\text{tor}}$ is non-trivial. As a working example we can take the CY $X^{\circ}$ of the previous section with $h^{1,1}(X^{\circ})=21$ and $h^{1,2}(X^{\circ})=1$, for which $B(X^{\circ})={\mathbb{Z}}_{5}$. The computation of $B(X^{\circ})$ was carried out in [1] as part of a study of integral cohomology of CY $3$-fold hypersurfaces in toric varieties. [^21] In that reference it was observed that hypersurfaces with non-trivial torsion subgroups are quite sparse, and in the enormous list of 473 800 776473800776473\,800\,776473 800 776 families there are just $32$ examples with torsion subgroups. Moreover, these come in $16$ mirror pairs, where each pair $(X,X^{\circ})$ consists of a manifold with $A(X)\neq 0$ and $B(X)=0$, while $A(X^{\circ})=0$ and $B(X^{\circ})\simeq A(X)$. The examples of section 2.1 show that this kind of simple relationship does not hold in general for CY $3$-folds. We saw that $B(X^{\circ})\neq 0$ means that the theory can be deformed by turning on a non-trivial gerbe, a choice of $\beta_{\text{t}}\in B(X^{\circ})$. The intuition from the path integral suggests that in the large radius limit we might think of a number of distinct complexified Kähler cones with continuous parameters $\beta_{\text{f}}+iJ\in H^{2}(X^{\circ},{\mathbb{C}})$ and distinguished by $\beta_{\text{t}}$ [25]. This leads to a number of questions, including:
@@ 4.1 Greene-Plesser mirrors via Landau-Ginzburg orbifolds
Rather than working directly with the abstract SCFT, we find it convenient to phrase the construction in terms of a Landau-Ginzburg orbifold (LGO) by using the correspondence between the $A_{d-2}$ minimal model and a (2,2) Landau-Ginzburg theory [82, 83] with a single chiral superfield $X$ and superpotential interaction $W=X^{d}$. This theory is believed to flow to the ${\cal A}_{d}=\text{A}_{d-2}$ minimal model [^23] We denote the $\text{A}_{d-2}$ model by ${\cal A}_{d}$ to avoid the $-2$ offset in the usual minimal model notation. with central charges $c=\overline{c}=3(1-2/d)$, and it has a symmetry $G={\mathbb{Z}}_{d}$, which acts non-chirally in the NS-NS sector by ${\boldsymbol{a}}\cdot X=e^{2\pi ia/d}X$. The SCFT has a trivial (a,c) ring, while the (c,c) ring is isomorphic to the quotient ring ${\cal R}={\mathbb{C}}[X]/\langle X^{d-1}\rangle$  determined by the superpotential.
@@
and whenever $F$ obeys the inclusions $F_{1}\subseteq F\subseteq G$, $F^{\circ}$ also satisfies the same inclusions. Thus, the orbifolds by phase symmetries that lead to (2,2) SCFTs with integral $\operatorname{U{}}(1)_{\text{L}}\times\operatorname{U{}}(1)_{\text{R}}$ charges come in pairs $F,F^{\circ}$, and mirror symmetry is the isomorphism ${\cal A}_{d}/F^{\circ}\simeq{\cal A}_{d}/F$.
@@ 4.2 Quintic quotients
We then see that ${\cal N}(F)/F\simeq K(F)\ltimes G_{0}/F$ We determined $K(F)$ for each of the quotients in the table, and we verified that, as might expected by mirror symmetry, $K(F)\simeq K(F^{\circ})$.
@@
where $\ell_{\alpha},m_{\alpha}\in{\mathbb{Z}}/5{\mathbb{Z}}$ The discrete torsion phase is parameterized by the integer $p$:
@@ A Gepner model for $X$ with $B(X)={\mathbb{Z}}_{3}$
for $\mathbf{b}=(b_{1},b_{2},b_{3}),~{}\mathbf{a}=(a_{1},a_{2},a_{3})\in F={\mathbb{Z}}_{9}^{R}\times{\mathbb{Z}}_{3}\times{\mathbb{Z}}_{9}$
@@
Although in this case the ambient toric orbifold has fixed surfaces, the example is rather like the quintic quotient ${\mathbb{P}}^{4}/F^{\circ}_{3}$: the choice of discrete torsion appears to be in one-to-one correspondence with $B(X)=G_{\text{dt}}={\mathbb{Z}}_{3}$, and the Hodge numbers do not jump because the twisted sectors that contain the marginal deformations that resolve the fixed surfaces are not affected by the discrete torsion. The next few examples display more elaborate possibilities, but all of them remain consistent with our proposal laid out above. [^28] Note that the phase ambiguities in a Landau-Ginzburg orbifold include an additional factor beyond the choice of discrete torsion. This factor arises as mod 2 ambiguity in the $G$–action on the Ramond-Ramond sector and is related to the Arf invariant and worldsheet spin structures [52, 85, 89]. When the projection to integral charges has an even order, this additional phase can lead to new possibilities of (2,2) orbifolds. In this work we set this phase to $+1$ and do not consider it further. It would be interesting to explore its implications for mirror symmetry and stringy geometry in detail.
@@ A Gepner model for $B(X)={\mathbb{Z}}_{2}$
Our examples are consistent with the inclusion $B(X)\subseteq G_{\text{dt}}$, and the jump in the Hodge numbers for $p\not\in B(X)$ is correlated with resolution of surface singularities in the ambient toric variety. We also note that for all of these examples we have known mirrors when $p=0$—these are manifolds with $\pi_{1}(X^{\circ})\simeq B(X)$ described in [1]. When $p\neq 0$, it is straightforward to extend the construction presented in section 4.2 to find explicit mirror SCFTs.

### v3

@@ 4 Flat gerbes and discrete torsion
In this section we turn to the situation where $X$ is simply connected but $B(X)=\{H^{3}(X,{\mathbb{Z}})\}_{\text{tor}}$ is non-trivial. As a working example we can take the CY $X^{\circ}$ of the previous section with $h^{1,1}(X^{\circ})=21$ and $h^{1,2}(X^{\circ})=1$, for which $B(X^{\circ})={\mathbb{Z}}_{5}$. The computation of $B(X^{\circ})$ was carried out in [1] as part of a study of integral cohomology of CY $3$-fold hypersurfaces in toric varieties. [^17] In that reference it was observed that hypersurfaces with non-trivial torsion subgroups are quite sparse, and in the enormous list of 473 800 776473\,800\,776 families there are just $32$ examples with torsion subgroups. Moreover, these come in $16$ mirror pairs, where each pair $(X,X^{\circ})$ consists of a manifold with $A(X)\neq 0$ and $B(X)=0$, while $A(X^{\circ})=0$ and $B(X^{\circ})\simeq A(X)$. The examples of section 2.1 show that this kind of simple relationship does not hold in general for CY $3$-folds. We saw that $B(X^{\circ})\neq 0$ means that the theory can be deformed by turning on a non-trivial gerbe, a choice of $\beta_{\text{t}}\in B(X^{\circ})$. The intuition from the path integral suggests that in the large radius limit we might think of a number of distinct complexified Kähler cones with continuous parameters $\beta_{\text{f}}+iJ\in H^{2}(X^{\circ},{\mathbb{C}})$ and distinguished by $\beta_{\text{t}}$ [25]. This leads to a number of questions, including:
@@ 4.1 Greene-Plesser mirrors via Landau-Ginzburg orbifolds
Rather than working directly with the abstract SCFT, we find it convenient to phrase the construction in terms of a Landau-Ginzburg orbifold (LGO) by using the correspondence between the $A_{d-2}$ minimal model and a (2,2) Landau-Ginzburg theory [82, 83] with a single chiral superfield $X$ and superpotential interaction $W=X^{d}$. This theory is believed to flow to the ${\cal A}_{d}=\text{A}_{d-2}$ minimal model [^19] We denote the $\text{A}_{d-2}$ model by ${\cal A}_{d}$ to avoid the 2-2 offset in the usual minimal model notation. with central charges $c=\overline{c}=3(1-2/d)$, and it has a symmetry $G={\mathbb{Z}}_{d}$, which acts non-chirally in the NS-NS sector by ${\boldsymbol{a}}\cdot X=e^{2\pi ia/d}X$. The SCFT has a trivial (a,c) ring, while the (c,c) ring is isomorphic to the quotient ring ${\cal R}={\mathbb{C}}[X]/\langle X^{d-1}\rangle$  determined by the superpotential.
@@
and whenever $F$ obeys the inclusions $_{1}\subseteq F\subseteq G, FF^{\circ}$ also satisfies the same inclusions. Thus, the orbifolds by phase symmetries that lead to (2,2) SCFTs with integral $\operatorname{U{}}(1)_{\text{L}}\times\operatorname{U{}}(1)_{\text{R}}$ charges come in pairs $F,F^{\circ}$, and mirror symmetry is the isomorphism ${\cal A}_{d}/F^{\circ}\simeq{\cal A}_{d}/F$.
@@ 4.2 Quintic quotients
We then see that ${\cal N}(F)/F\simeq K(F)\ltimes G_{0}/F$ . We determined $K(F)$ for each of the quotients in the table, and we verified that, as might expected by mirror symmetry, $K(F)\simeq K(F^{\circ})$.
@@
where $\ell_{\alpha},m_{\alpha}\in{\mathbb{Z}}/5{\mathbb{Z}}$ . The discrete torsion phase is parameterized by the integer $p$:
@@ A Gepner model for $X$ with $B(X)={\mathbb{Z}}_{3}$
for $\mathbf{b}=(b_{1},b_{2},b_{3}),~\mathbf{a}=(a_{1},a_{2},a_{3})\in F={\mathbb{Z}}_{9}^{R}\times{\mathbb{Z}}_{3}\times{\mathbb{Z}}_{9}$ .
@@
Although in this case the ambient toric orbifold has fixed surfaces, the example is rather like the quintic quotient ${\mathbb{P}}^{4}/F^{\circ}_{3}$: the choice of discrete torsion appears to be in one-to-one correspondence with $B(X)=G_{\text{dt}}={\mathbb{Z}}_{3}$, and the Hodge numbers do not jump because the twisted sectors that contain the marginal deformations that resolve the fixed surfaces are not affected by the discrete torsion. The next few examples display more elaborate possibilities, but all of them remain consistent with our proposal laid out above. [^24] Note that the phase ambiguities in a Landau-Ginzburg orbifold include an additional factor beyond the choice of discrete torsion. This factor arises as mod 2 ambiguity in the GG–action on the Ramond-Ramond sector and is related to the Arf invariant and worldsheet spin structures [52, 85, 89]. When the projection to integral charges has an even order, this additional phase can lead to new possibilities of (2,2) orbifolds. In this work we set this phase to +1+1 and do not consider it further. It would be interesting to explore its implications for mirror symmetry and stringy geometry in detail.
@@ A Gepner model for $B(X)={\mathbb{Z}}_{2}$
Our examples are consistent with the inclusion $B(X)\subseteq G_{\text{dt}}$, and the jump in the Hodge numbers for $p\not\in B(X)$ is correlated with resolution of surface singularities in the ambient toric variety. We also note that for all of these examples we have known mirrors when p=0p=0—these are manifolds with $\pi_{1}(X^{\circ})\simeq B(X)$ described in [1]. When $p\neq 0$, it is straightforward to extend the construction presented in section 4.2 to find explicit mirror SCFTs.